a partical droped from the heighest place of a building. in the final second the particle moved half height of the bulding . what is the height of the building?

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago

distance travelled in n seconds, s = 5n^2
v= 10(n-1)
s’= 10(n-1)+ 5 (n-1)^2
s-s’= s/2= 5n^2- 10n+10-5n^2+10n-5= 5= s/2
s=10